OpenStudy (anonymous):

find the height of the triangle A. 83.1 B.46.1 C.171.4 D.108.6

OpenStudy (anonymous):



OpenStudy (anonymous):

use pythagoras theorem.... http://www.mathsisfun.com/pythagoras.html does that help??

OpenStudy (anonymous):

not really cause i only have the length of one side. but thanks anyway

OpenStudy (anonymous):

oh sorry.....you don't need to use that

OpenStudy (anonymous):

\[\sin(61)=\frac{x}{95}\iff x=95\times \sin(61)\]

OpenStudy (anonymous):

alright so the answers A then. thank you!

OpenStudy (anonymous):

yw